(x-h)^2=x^2-8x+14
Find the value of h
1
Expert's answer
2019-05-21T09:04:57-0400

"(x-h)^2=x^2-8x+14"

"h^2-2xh+8x-14=0"




"h_1=\sqrt{\smash[b]{x^{2}-8x+14}}+x""h_2=x-\sqrt{\smash[b]{x^{2}-8x+14}}"